# Rotatable Antenna Enabled Wireless Communication: Modeling and Optimization

### v2

@@ VI Conclusion
In this paper, we proposed a new RA model that provides new spatial DoFs to reconfigure the array directional gain pattern by flexibly adjusting the 3D orientation/boresight of each antenna, thus substantially enhancing the array gain and transmission rate without increasing the number of antennas or changing antenna positions. Specifically, the receive beamforming and the deflection angles of all RAs were jointly optimized to maximize the minimum SINR among all users. The optimal closed-form RA deflection angles were first derived with MRC receive beamforming applied in the single-user and free-space propagation setup. Meanwhile, the asymptotic performance analysis for the case with an infinitely large number of antennas showed that the RA-enabled system always achieves a higher array gain as compared to the conventional fixed-antenna system. For the general multi-user and multi-path channel setup, an AO algorithm and a two-stage algorithm were proposed to obtain high-quality suboptimal solutions to balance the array directional gain over the multi-path channels. Simulation results validated our analytical results and demonstrated that our proposed RA-enabled system significantly outperforms various benchmark schemes. It was shown that even with a small deflection angle range for RA orientation/boresight adjustment, the RA-enabled system could still reap considerable performance gains.
@@ Appendix A Proof of Theorem 1
First, for the case of $p=\frac{1}{2}$ and $N_{x}\leq\hat{N}_{x}\triangleq 2\left\lfloor\frac{\operatorname{tan}\theta_{\mathrm{max}}}{\delta}\right\rfloor+1$, the integral in (22) can be simplified as
$\mathbb{D}=$ $\int_{-\frac{N_{x}d}{2}}^{\frac{N_{x}d}{2}}\frac{1}{1+\frac{x^{2}}{r^{2}}}\mathrm{d}x=2r\int_{0}^{\frac{N_{x}\delta}{2}}{\frac{1}{1+x^{2}}\mathrm{d}x}$
$=$ $2r\left.\operatorname{arctan}x\right|_{0}^{\frac{\hat{N}_{x}\delta}{2}}=2r\operatorname{arctan}\frac{\hat{N}_{x}\delta}{2}.$ (52)
Then, for $N_{x}>\hat{N}_{x}$, the integral in (22) can be calculated as
$\mathbb{D}=$ $2\left[\int_{0}^{\frac{\hat{N}_{x}d}{2}}{\frac{1}{1+\frac{x^{2}}{r^{2}}}\mathrm{d}x}+\int_{\frac{\hat{N}_{x}d}{2}}^{\frac{N_{x}d}{2}}{\frac{\operatorname{cos}\left(\operatorname{arctan}|\frac{x}{r}|-\theta_{\mathrm{max}}\right)}{1+\frac{x^{2}}{r^{2}}}\mathrm{d}x}\right]$
=$\overset{(a)}{=}$ $2r\left[\int_{0}^{\frac{\hat{N}_{x}\delta}{2}}{\frac{1}{1+x^{2}}\mathrm{d}x}+\int_{\frac{\hat{N}_{x}\delta}{2}}^{\frac{N_{x}\delta}{2}}{\frac{\operatorname{cos}\theta_{\mathrm{max}}+\operatorname{sin}\theta_{\mathrm{max}}x}{(1+x^{2})^{\frac{3}{2}}}\mathrm{d}x}\right]$
=$\overset{(b)}{=}$ $2r\left[\left.\operatorname{arctan}x\right|_{0}^{\frac{\hat{N}_{x}\delta}{2}}+\left.\frac{\operatorname{cos}\theta_{\mathrm{max}}x-\operatorname{sin}\theta_{\mathrm{max}}}{\sqrt{1+x^{2}}}\right|_{\frac{\hat{N}_{x}\delta}{2}}^{\frac{N_{x}\delta}{2}}\right]$
=$\overset{(c)}{=}$ $2r\left[\theta_{\mathrm{max}}+\operatorname{sin}\left(\operatorname{arctan}\frac{N_{x}\delta}{2}-\theta_{\mathrm{max}}\right)\right],$ (53)
where $(a)$ holds due to $\operatorname{cos}(\operatorname{arctan}(x))=\frac{1}{\sqrt{1+x^{2}}}$ and $\operatorname{sin}(\operatorname{arctan}(x))=\frac{x}{\sqrt{1+x^{2}}}$, $(b)$ follows the integral formulas 2.103.4, 2.264.5 and 2.264.6 in [34], and $(c)$ holds due to the fact that $\operatorname{arctan}\frac{\hat{N}_{x}\delta}{2}=\theta_{\mathrm{max}}$.
@@ Appendix B Proof of Lemma 2
For the case of $\sqrt{(N_{x}d)^{2}+(N_{y}d)^{2}}\leq 2r\operatorname{tan}\theta_{\mathrm{max}}$, i.e., $\epsilon_{n}=0,\;\forall n$, the SNR expression in (21) reduces to
$\gamma=\frac{\bar{P}G_{0}\xi\delta^{2}}{4\pi d^{2}}\int_{-\frac{N_{y}d}{2}}^{\frac{N_{y}d}{2}}{\int_{-\frac{N_{x}d}{2}}^{\frac{N_{x}d}{2}}{\frac{1}{1+\frac{1}{r^{2}}(x^{2}+y^{2})}\mathrm{d}x\mathrm{d}y}}.$ (54)
@@
=$\overset{(d)}{=}$ $r^{2}\int_{-\frac{N_{y}\delta}{2}}^{\frac{N_{y}\delta}{2}}\left.\frac{1}{\sqrt{1+y^{2}}}\operatorname{arctan}\frac{x}{\sqrt{1+y^{2}}}\mathrm{d}y\right|_{-\frac{N_{x}\delta}{2}}^{\frac{N_{x}\delta}{2}}$
$=$ $r^{2}\int_{-\frac{N_{y}\delta}{2}}^{\frac{N_{y}\delta}{2}}\frac{2}{\sqrt{1+y^{2}}}\operatorname{arctan}\frac{N_{x}\delta}{2\sqrt{1+y^{2}}}\mathrm{d}y$
@@
$=$ $\left.\frac{1}{4}N_{x}\pi\delta r^{2}\operatorname{arctan}y\right|_{-\frac{N_{y}\delta}{2}}^{\frac{N_{y}\delta}{2}}$
$=$ $\frac{1}{2}N_{x}\pi\delta r^{2}\operatorname{arctan}\frac{N_{y}\delta}{2}$
@@ Appendix C Proof of Theorem 2
$\mathcal{F}(R,p,\theta_{\mathrm{max}})=\frac{\bar{P}G_{0}\xi\delta^{2}}{4\pi d^{2}}\left(\int_{0}^{2\pi}\mathrm{d}\zeta\int_{0}^{D}\frac{1}{1+\left(\frac{l}{r}\right)^{2}}l\mathrm{d}l+\right.$
$\quad\quad\quad\quad\left.\int_{0}^{2\pi}\mathrm{d}\zeta\int_{D}^{R}\frac{\operatorname{cos}^{2p}\left(\operatorname{arctan}\frac{l}{r}-\theta_{\mathrm{max}}\right)}{1+\left(\frac{l}{r}\right)^{2}}l\mathrm{d}l\right),$ (57)
@@
$\mathbb{D}_{1}=$ $2\pi r^{2}\int_{0}^{\frac{D}{r}}\frac{l\mathrm{d}l}{1+l^{2}}=\pi r^{2}\left.\operatorname{ln}(1+l^{2})\right|_{0}^{\frac{D}{r}}$
@@
$\mathbb{D}_{2}=$ $2\pi r^{2}\int_{\frac{D}{r}}^{\frac{R}{r}}\frac{\operatorname{cos}^{2p}\left(\operatorname{arctan}l-\theta_{\mathrm{max}}\right)}{1+l^{2}}l\mathrm{d}l$
$=$ $2\pi r^{2}\int_{\operatorname{arctan}\frac{D}{r}-\theta_{\mathrm{max}}}^{\operatorname{arctan}\frac{R}{r}-\theta_{\mathrm{max}}}\operatorname{cos}^{2p}\epsilon\operatorname{tan}(\epsilon+\theta_{\mathrm{max}})\mathrm{d}\epsilon.$ (59)
@@ Appendix D Proof of Lemma 3
$\mathbb{D}=$ $\int_{\operatorname{arctan}\frac{D}{r}}^{\operatorname{arctan}\frac{R}{r}}\operatorname{cos}(\epsilon-\theta_{\mathrm{max}})\operatorname{tan}\epsilon\mathrm{d}\epsilon$
$=$ $\int_{\operatorname{arctan}\frac{D}{r}}^{\operatorname{arctan}\frac{R}{r}}\left(\operatorname{cos}\theta_{\mathrm{max}}\operatorname{sin}\epsilon+\operatorname{sin}\theta_{\mathrm{max}}\frac{\operatorname{sin}^{2}\epsilon}{\operatorname{cos}\epsilon}\right)\mathrm{d}\epsilon$
$=$ $\operatorname{cos}\theta_{\mathrm{max}}\left(\operatorname{cos}\left(\operatorname{arctan}\frac{D}{r}\right)-\operatorname{cos}\left(\operatorname{arctan}\frac{R}{r}\right)\right)+$
$\quad\operatorname{sin}\theta_{\mathrm{max}}\int_{\operatorname{sin}\left(\operatorname{arctan}\frac{D}{r}\right)}^{\operatorname{sin}\left(\operatorname{arctan}\frac{R}{r}\right)}\frac{\epsilon^{2}}{1-\epsilon^{2}}\mathrm{d}\epsilon$
@@
where $\theta_{R}\triangleq\operatorname{arctan}\frac{R}{r}$, $(g)$ follows the integral formulas 2.147.1 and 2.111.6 in [34], and it has $\theta_{\mathrm{max}}=\operatorname{arctan}\frac{D}{r}$ when $R\geq D$.

### v4

@@ VI Conclusion
In this paper, we proposed a new RA model that provides new spatial DoFs to reconfigure the array directional gain pattern by flexibly adjusting the 3D orientation/boresight of each antenna, thus substantially enhancing the array gain and transmission rate without increasing the number of antennas or changing antenna positions. Specifically, the receive beamforming and the pointing vectors of all RAs were jointly optimized to maximize the minimum SINR among all users. The optimal closed-form pointing vectors for all RAs were first derived with MRC receive beamforming applied in the single-user and free-space propagation setup. Meanwhile, the asymptotic performance analysis for the case with an infinite number of antennas demonstrated that the RA system consistently achieves a higher array gain compared to the conventional fixed-antenna system. For the general multi-user and multipath channel setup, an AO algorithm and a two-stage algorithm were proposed to obtain high-quality suboptimal solutions to balance the array directional gain over the multipath channels. Simulation results validated our analytical results and demonstrated that our proposed RA system significantly outperforms various benchmark schemes. It was shown that even with a small rotational range for RA orientation/boresight adjustment, the RA system could still reap considerable performance gains.
@@ Appendix A Proof of Theorem 1
First, for the case of $p=\frac{1}{2}$ and $N_{x}\leq\bar{N}_{x}\triangleq 2\left\lfloor\frac{\operatorname{tan}\theta_{\mathrm{max}}}{\delta}\right\rfloor+1$, the integral in (25) can be simplified as
$\mathbb{D}=$ $\int_{-\frac{N_{x}\Delta}{2}}^{\frac{N_{x}\Delta}{2}}\frac{1}{1+\frac{x^{2}}{r^{2}}}\mathrm{d}x=2r\int_{0}^{\frac{N_{x}\delta}{2}}{\frac{1}{1+x^{2}}\mathrm{d}x}$
$=$ $2r\left.\operatorname{arctan}\left(x\right)\right|_{0}^{\frac{\bar{N}_{x}\delta}{2}}=2r\operatorname{arctan}\left(\frac{\bar{N}_{x}\delta}{2}\right).$ (56)
Then, for $N_{x}>\bar{N}_{x}$, the integral in (25) can be calculated as
$\mathbb{D}=$ $2\left[\int_{0}^{\frac{\bar{N}_{x}\Delta}{2}}{\frac{1}{1+\frac{x^{2}}{r^{2}}}\mathrm{d}x}+\int_{\frac{\bar{N}_{x}\Delta}{2}}^{\frac{N_{x}\Delta}{2}}{\frac{\operatorname{cos}\left(\operatorname{arctan}\left(|\frac{x}{r}|\right)-\theta_{\mathrm{max}}\right)}{1+\frac{x^{2}}{r^{2}}}\mathrm{d}x}\right]$
=$\overset{(a)}{=}$ $2r\left[\int_{0}^{\frac{\bar{N}_{x}\delta}{2}}{\frac{1}{1+x^{2}}\mathrm{d}x}+\int_{\frac{\bar{N}_{x}\delta}{2}}^{\frac{N_{x}\delta}{2}}{\frac{\operatorname{cos}\theta_{\mathrm{max}}+\operatorname{sin}\theta_{\mathrm{max}}x}{(1+x^{2})^{\frac{3}{2}}}\mathrm{d}x}\right]$
=$\overset{(b)}{=}$ $2r\left[\left.\operatorname{arctan}\left(x\right)\right|_{0}^{\frac{\bar{N}_{x}\delta}{2}}+\left.\frac{\operatorname{cos}\theta_{\mathrm{max}}x-\operatorname{sin}\theta_{\mathrm{max}}}{\sqrt{1+x^{2}}}\right|_{\frac{\bar{N}_{x}\delta}{2}}^{\frac{N_{x}\delta}{2}}\right]$
=$\overset{(c)}{=}$ $2r\left[\theta_{\mathrm{max}}+\operatorname{sin}\left(\operatorname{arctan}\left(\frac{N_{x}\delta}{2}\right)-\theta_{\mathrm{max}}\right)\right],$ (57)
where $(a)$ holds due to $\operatorname{cos}(\operatorname{arctan}(x))=\frac{1}{\sqrt{1+x^{2}}}$ and $\operatorname{sin}(\operatorname{arctan}(x))=\frac{x}{\sqrt{1+x^{2}}}$, $(b)$ follows the integral formulas 2.103.4, 2.264.5, and 2.264.6 in [47], and $(c)$ holds due to the fact that $\operatorname{arctan}\left(\frac{\bar{N}_{x}\delta}{2}\right)=\theta_{\mathrm{max}}$.
@@ Appendix B Proof of Lemma 2
For the case of $\sqrt{(N_{x}\Delta)^{2}+(N_{y}\Delta)^{2}}\leq 2r\operatorname{tan}\theta_{\mathrm{max}}$, i.e., $\epsilon_{n}=0,\;\forall n$, the SNR expression in (24) reduces to
$\gamma=\frac{\bar{P}G_{0}\xi\delta^{2}}{4\pi\Delta^{2}}\int_{-\frac{N_{y}\Delta}{2}}^{\frac{N_{y}\Delta}{2}}{\int_{-\frac{N_{x}\Delta}{2}}^{\frac{N_{x}\Delta}{2}}{\frac{1}{1+\frac{1}{r^{2}}(x^{2}+y^{2})}\mathrm{d}x\mathrm{d}y}}.$ (58)
@@
=$\overset{(d)}{=}$ $r^{2}\int_{-\frac{N_{y}\delta}{2}}^{\frac{N_{y}\delta}{2}}\left.\frac{1}{\sqrt{1+y^{2}}}\operatorname{arctan}\left(\frac{x}{\sqrt{1+y^{2}}}\right)\mathrm{d}y\right|_{-\frac{N_{x}\delta}{2}}^{\frac{N_{x}\delta}{2}}$
$=$ $r^{2}\int_{-\frac{N_{y}\delta}{2}}^{\frac{N_{y}\delta}{2}}\frac{2}{\sqrt{1+y^{2}}}\operatorname{arctan}\left(\frac{N_{x}\delta}{2\sqrt{1+y^{2}}}\right)\mathrm{d}y$
@@
$=$ $\left.\frac{1}{4}N_{x}\pi\delta r^{2}\operatorname{arctan}\left(y\right)\right|_{-\frac{N_{y}\delta}{2}}^{\frac{N_{y}\delta}{2}}$
$=$ $\frac{1}{2}N_{x}\pi\delta r^{2}\operatorname{arctan}\left(\frac{N_{y}\delta}{2}\right)$
@@ Appendix C Proof of Theorem 2
$\mathcal{F}(R,p,\theta_{\mathrm{max}})=\frac{\bar{P}G_{0}\xi\delta^{2}}{4\pi\Delta^{2}}\left(\int_{0}^{2\pi}\mathrm{d}\zeta\int_{0}^{D}\frac{1}{1+\left(\frac{l}{r}\right)^{2}}l\mathrm{d}l+\right.$
$\quad\quad\quad\quad\left.\int_{0}^{2\pi}\mathrm{d}\zeta\int_{D}^{R}\frac{\operatorname{cos}^{2p}\left(\operatorname{arctan}\left(\frac{l}{r}\right)-\theta_{\mathrm{max}}\right)}{1+\left(\frac{l}{r}\right)^{2}}l\mathrm{d}l\right),$ (61)
@@
$\mathbb{D}_{1}=$ $2\pi r^{2}\int_{0}^{\frac{D}{r}}\frac{l}{1+l^{2}}\mathrm{d}l=\pi r^{2}\left.\operatorname{ln}(1+l^{2})\right|_{0}^{\frac{D}{r}}$
@@
$\mathbb{D}_{2}=$ $2\pi r^{2}\int_{\frac{D}{r}}^{\frac{R}{r}}\frac{\operatorname{cos}^{2p}\left(\operatorname{arctan}\left(l\right)-\theta_{\mathrm{max}}\right)}{1+l^{2}}l\mathrm{d}l$
$=$ $2\pi r^{2}\int_{\operatorname{arctan}\left(\frac{D}{r}\right)-\theta_{\mathrm{max}}}^{\operatorname{arctan}\left(\frac{R}{r}\right)-\theta_{\mathrm{max}}}\operatorname{cos}^{2p}\epsilon\operatorname{tan}(\epsilon+\theta_{\mathrm{max}})\mathrm{d}\epsilon.$ (63)
@@ Appendix D Proof of Lemma 3
$\mathbb{D}=$ $\int_{\operatorname{arctan}\left(\frac{D}{r}\right)}^{\operatorname{arctan}\left(\frac{R}{r}\right)}\operatorname{cos}(\epsilon-\theta_{\mathrm{max}})\operatorname{tan}\epsilon\mathrm{d}\epsilon$
$=$ $\int_{\operatorname{arctan}\left(\frac{D}{r}\right)}^{\operatorname{arctan}\left(\frac{R}{r}\right)}\left(\operatorname{cos}\theta_{\mathrm{max}}\operatorname{sin}\epsilon+\operatorname{sin}\theta_{\mathrm{max}}\frac{\operatorname{sin}^{2}\epsilon}{\operatorname{cos}\epsilon}\right)\mathrm{d}\epsilon$
$=$ $\operatorname{cos}\theta_{\mathrm{max}}\left(\operatorname{cos}\left(\operatorname{arctan}\left(\frac{D}{r}\right)\right)-\operatorname{cos}\left(\operatorname{arctan}\left(\frac{R}{r}\right)\right)\right)+$
$\quad\operatorname{sin}\theta_{\mathrm{max}}\int_{\operatorname{sin}\left(\operatorname{arctan}\left(\frac{D}{r}\right)\right)}^{\operatorname{sin}\left(\operatorname{arctan}\left(\frac{R}{r}\right)\right)}\frac{\epsilon^{2}}{1-\epsilon^{2}}\mathrm{d}\epsilon$
@@
where $\theta_{R}\triangleq\operatorname{arctan}\left(\frac{R}{r}\right)$, $(g)$ follows the integral formulas 2.147.1 and 2.111.6 in [47], and it has $\theta_{\mathrm{max}}=\operatorname{arctan}\left(\frac{D}{r}\right)$ when $R\geq D$.